OpenStudy (lizeth):

(x-6)(x-3i)(x+3i) in standard form

OpenStudy (anonymous):

(x-6)(x-3i)(x+3i) leave the first term and first of all multiply next two terms by formula that is (a+b)(a-b) = a^2 - b^2

hero (hero):

(x-3i)(x+3i) = (x^2 - 9i^2) (x-6)(x^2 - 9i^2) = x^2 - 9i^2(x-6) = x^2(x-6) - 9i^2(x-6) = x^3 - 6x^2 - 9xi^2 + 54i^2 = x^3 - 6x^2 + 45i^2